OpenStudy (anonymous):

of sin theta =2/3 find cos theta show work i need help asap edit-hint sin 2 theta+ cos2 theta = 0

Nnesha (nnesha):

 you should be familiar with\[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\]

Nnesha (nnesha):

so draw a right triangle given is sin theta =2/3 so what would be opposite side of theta ?>

OpenStudy (anonymous):

idk tbh im not good at this

OpenStudy (anonymous):

adjacent?

Nnesha (nnesha):

\[\huge\rm \sin \theta =\frac{\color{REd}{opposite}}{\color{blue}{hypotenuse}}=\frac{\color{red}{2}}{\color{blue}{3}}\]

OpenStudy (anonymous):

and then

Nnesha (nnesha):

numerator = length of opposite side denominator = length of hypohtunse side

Nnesha (nnesha):

i just saw the hint you can use that as well substitue sin for 2/3 and then solve for cos

OpenStudy (anonymous):

can you write out the steps

OpenStudy (anonymous):

to all i need t see it

Nnesha (nnesha):

the hint should be this \[\huge\rm \sin^2 \theta + \cos^2 \theta =1\] equal to 1 not 0 ?

OpenStudy (anonymous):

oh my bad

OpenStudy (anonymous):

you are right sorry lol

Nnesha (nnesha):

its okay that's the MOST famous identity alright you can use that equation to find cos substitute sin for 2/3

OpenStudy (anonymous):

so 2/3+cos theta=1

Nnesha (nnesha):

\[\huge\rm \sin^2 \theta + \cos^2 \theta =1\] \[\huge\rm (\frac{2}{3})^2 + \cos^2 \theta =1\] solve for cos

Nnesha (nnesha):

don't forget the squares

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

help me solve?